Buttercup Inc. just issued RM1,000 par 30-year bonds. Each bond was sold for
RM1,107.20 and pay interest semiannually. Investors require a rate of 7.75% on the
bonds. What is the bonds' coupon rate?
1
Expert's answer
2015-03-17T13:03:26-0400
RM1,000 par 30-year bonds. Each bond was sold for RM1,107.20 and pay interest semiannually. Required rate of return 7.75%.
A coupon payment on a bond is a periodic interest payment that the
bondholder receives during the time between when the bond is issued and
when it matures.
For bonds:
P = F*if((1 - (1 + i)^-n)/i) + M(1 + i)^-n
where:
C = F * iF = coupon payment
N = number of payments
i = market interest rate, or required yield, M = face value
P = market price of bond.
The bonds' coupon rate if = (P - M(1 + i)^-n)/(F*((1 - (1 + i)^-n)/i)) =
(1107.2 - 1000/1.0775^30)/(1000*((1 - 1/1.0775^30)/0.0775)) = 8.68%.